# A Neural Difference-of-Entropies Estimator for Mutual Information

### v1

@@ Abstract
Estimating Mutual Information (MI), a key measure of dependence of random quantities without specific modelling assumptions, is a challenging problem in high dimensions. We propose a novel mutual information estimator based on parametrizing conditional densities using normalizing flows, a deep generative model that has gained popularity in recent years. This estimator leverages a block autoregressive structure to achieve improved bias-variance trade-offs on standard benchmark tasks.
@@ 1 Introduction
Mutual Information (MI), a measure of dependence of random variables $X$ and $Y$, plays an important role in information theory (Cover & Thomas, 2006), statistics and machine learning (Tishby et al., 2000; Peng et al., 2005; Vergara & Estévez, 2014; Chen et al., 2018), and biology and medical sciences (Zhang et al., 2012; Sengupta et al., 2022). For random variables $X$ and $Y$ with joint density $p$, the mutual information is defined as
@@
Formally, an MI estimator $\hat{I}_{N}$ depends on independent and identically distributed (i.i.d.) random sample pairs $(X_{1},Y_{1}),\dots,(X_{N},Y_{N})$, and should ideally be unbiased, consistent and efficient. In addition, such an estimator should be effectively computable from large and high-dimensional data. We propose an unbiased and consistent mutual information estimator based on the difference-of-entropies (DoE) estimator suggested in McAllester & Stratos (2018). This characterization expresses the mutual information as the difference between the entropy of $X$ and the conditional entropy of $X$ given $Y$ ,
$I(X;Y)=H(X)-H(X\ |\ Y).$
@@ 1.1 Overview of previous work
Traditional MI estimators are non-parametric estimators that depend on density estimation and Monte Carlo integration or on the computation of $k$ nearest neighbours (kNN). Examples include the widely used KSG estimator by Kraskov et al. (2004), the non-parametric kNN estimator (kpN) by Lombardi & Pant (2015), and improvements of the KSG estimator and a geometric kNN estimator by Gao et al. (2017). These non-parametric methods are fast and accurate for low-dimensional and small-sized problems and are easy to implement. However, they suffer from the curse of dimensionality and do not scale well in machine learning problems since the data sets can be relatively large and high-dimensional (Paninski, 2003).
@@
More recent parametric methods take advantage of deep learning architectures to approximate variational bounds on MI. These have been categorized into discriminative and generative approaches by Song & Ermon (2019). Some state-of-art discriminative approaches include InfoNCE (van den Oord et al., 2018), MINE (Belghazi et al., 2018), SMILE (Song & Ermon, 2019), CCMI (Mukherjee et al., 2019) and DEMI (Liao et al., 2020). van den Oord et al. (2018) proposed a contrastive predictive coding (CPC) method which relies on maximizing a lower bound on mutual information. The lower bound involves function approximators implemented with neural networks and is constrained by the batch size $N$, leading to a method that is more biased but with less variance. MINE, on the other hand, is based on the Donsker-Varadhan (DV) lower bound for the KL divergence. Fundamental limitations on approaches based on variational bounds were studied by  Song & Ermon (2019) and McAllester & Stratos (2018), the latter being the motivation for our approach.
@@
Instead of constructing the Mutual Information estimators based on variational lower bounds, Liao et al. (2020) proposed a classifier-based estimator called DEMI, where a parametrized classifier is trained to distinguish between the joint density $p(x,y)$ and the product $p(x)p(y)$. Mukherjee et al. (2019) proposed another classifier-based (conditional) MI estimator that is asymptotically equivalent to DEMI. However, it still relies on variational lower bounds and is prone to higher error than DEMI for finite samples, as summarized by Liao et al. (2020).
@@
Compared with discriminative approaches, generative approaches are less well explored in MI estimation problems. A naïve approach using generative models for estimating MI is to learn the entropies $H(X),H(Y)$ and $H(X,Y)$ with three individual generative models, such as VAE (Kingma & Welling, 2013) or Normalizing Flows (Kobyzev et al., 2020), from samples. A method for estimating entropy using normalizing flows was introduced by Ao & Li (2022). Estimators based on the individual entropies will be highly biased and computationally expensive since the entropies are trained separately, while it is revealed that considering the enhancement of correlation between entropies in constructing MI estimators can improve the bias (Gao et al., 2017). Duong & Nguyen (2023) propose the Diffeomorphic Information Neural Estimator (DINE), that takes advantage of the invariance of conditional mutual information under diffeomorphisms. An alternative approach to estimating MI using normalizing flows was recently proposed by Butakov et al. (2024). This approach takes advantage of the invariance of the pointwise mutual information under diffeomorphisms. In addition, the methods from Butakov et al. (2024) allow for the estimation of mutual information using direct, closed-form expressions. Finally, we would like to point the recently introduced MINDE estimator Franzese et al. (2024), which is based on diffusion models and represents a complementary approach.
@@
From a practical point of view, the performance of MI estimators is often measured using standard data sets based on Gaussian distributions for which the ground truth is known. Recently, Czyż et al. (2023) proposed a collection of benchmarks to evaluate the performance of different MI estimators in more challenging environments.
@@ 2.1 Mutual Information and Entropy
Given a pair of random variables $(X,Y)$, the conditional entropy of $X$ with respect to $Y$ is defined as $H(X|Y)=H(X,Y)-H(Y)$, where $H(X,Y)$ is the joint entropy of $(X,Y)$ (not to be confused with the cross-entropy, introduced below). The mutual information can be expressed in terms of entropies via the 3H principle:
@@
The characterization (*) is the basis of the difference-of-entropies (DoE) estimator introduced by McAllester & Stratos (2018).
@@
$Q(p,q):=-\mathbb{E}_{p}[\log q(X)].$
One easily checks that the cross-entropy, entropy and KL divergence are related via
@@
Let $(X,Y)$ be a pair of random variable with joint density $p$. Then
@@
Given data $\{(x_{i},y_{i})\}_{i=1}^{N}$, the resulting difference-of-entropies (DoE) estimator, as suggested by McAllester & Stratos (2018), consists of minimizing the objectives
$\hat{Q}(p_{X},q_{X})=-\frac{1}{N}\sum_{i=1}^{N}\log q_{X}(x_{i}),\quad\hat{Q}(p,q_{X|Y})=-\frac{1}{N}\sum_{i=1}^{N}\log q_{X|Y}(x_{i}|y_{i})$ (4)
@@ 2.2 Normalizing flows
[…]
@@
It is known that any continuous distribution with density $p_{X}$ satisfying some mild conditions can be generated from the uniform distribution on a cube $[0,1]^{n}$ (and hence, by invertibility, from any other distribution satisfying the same conditions) if the transformation $f$ can have arbitrary complexity (Bogachev et al., 2007). However, as is common with universal approximation results, this result does not translate into a practical recipe. A more practical approach is to use a composition of simple functions implemented by neural networks, which have sufficient expressive power. An obvious but important property of diffeomorphisms is that they are composable. Specifically, let $g_{1},g_{2},\dots,g_{K}$ be a set of $K$ diffeomorphisms and denote by $g=g_{K}\circ g_{K-1}\circ\cdots\circ g_{1}$ the composition of these functions. The determinant of the Jacobian is then given by
@@
In light of equation 2, minimizing the forward KL divergence is equivalent to minimizing the cross-entropy
$\begin{split}\mathcal{L}(\Theta)&:=Q(p_{X}^{*},p_{X}(\cdot;\Theta))=-\mathbb{E}_{p_{X}^{*}}[\log p_{X}(X;\Theta)]\\ &=-\mathbb{E}_{p_{X}^{*}}[\log p_{Z}(f_{\phi}(X);\psi)+\log|\det\textnormal{d}f_{\phi}(X)|].\end{split}$ (6)
@@
^$\hat{\mathcal{L}}(\Theta):=-\frac{1}{N}\sum_{j=1}^{N}(\log p_{Z}(f_{\phi}(x_{j});\psi)+\log|\det\textnormal{d}f_{\phi}(x_{j})|).$ (7)
@@ 2.2.1 Block autoregressive flows
Autoregressive flows (Kingma et al., 2016) are normalizing flows with the convenient property that their Jacobian is triangular. Block neural autoregressive flows (B-NAF), introduced by De Cao et al. (2019), are flows that are autoregressive and monotone, but that are implemented using a single neural network architecture, rather than relying on conditioner networks. More specifically, a block autoregressive flow is given as a sequence of transformations
@@
where each $B_{ij}^{(k)}\in R^{a_{k+1}\times a_{k}}$ and $g(x)=\exp(x)$ applied componentwise, to ensure that the entries are positive. We set $a_{0}=a_{\ell+1}=1$. It is not hard to see that the $i$-component of $f(x)$ only depends on $x_{1},\dots,x_{i}$. Since the product of block diagonal matrices with blocks of size $a\times b$ and $b\times c$, respectively, is block diagonal with size $a\times c$, the composition $f$ has lower triangular Jacobian with positive diagonal entries, and hence is invertible. The determinant of the triangular Jacobian matrix is the product of the diagonal entries $\partial f_{i}/\partial x_{i}$, each of which can be computed as product
@@ 3 Joint estimation of Mutual Information
The Jacobian determinant is the product of the partial derivative $\partial f_{1}/\partial y$ and $\partial f_{2}/\partial x$ (see Section 2.2.1). Suppose $p(x,y)$ is a standard Gaussian density, so that $p(x,y)=p_{X}(x)p_{Y}(y)$, and that we have data $(x_{i},y_{i})$ from an unknown distribution $q$. The cost function equation 7 for learning a normalizing flow takes the form
$-\frac{1}{N}\sum_{i=1}^{N}\left(\log p_{X}(f_{2}(y_{i},x_{i}))+\log\frac{\partial f_{2}}{\partial x}(y_{i},x_{i})\right)+\left(\log p_{Y}(f_{1}(y_{i}))+\log\frac{\partial f_{1}}{\partial y}(y_{i})\right).$ (8)
The components $f_{1}$ and $f_{2}$ depend on a distinct set of weights in the neural network. Optimizing only the part of equation 8 involving $f_{1}$ on data $\{y_{i}\}$ gives an estimate for the entropy of $Y$, while optimizing the part with $f_{2}$ on data $\{(x_{i},y_{i})\}$ gives rise to an estimate of the cross-entropy $H(X\ |\ Y)$. Moreover, if we deactivate the weights in off-diagonal blocks (the dashed lines), then optimizing this part on data $\{x_{i}\}$ gives an estimate of $H(X)$. Note that training for $H(X\ |\ Y)$ and setting the off-diagonal weights to zero does not automatically give an estimators for $H(X)$. It is, however, conceivable that one can begin with a network that approximates $H(X)$ and then optimize the off-diagonal weights to obtain an approximation of $H(X|Y)$. A justification of this approach is provided in the appendix.
@@
where now we simply write $p$ for the density of a Gaussian. Optimizing this function gives an estimate of the cross-entropy $H(X\ |\ Y)$. If, on the other hand, we set the off-diagonal weights to zero and optimize the resulting function $\mathcal{L}_{2}$, we get an estimator for the entropy $H(X)$. This motivates Algorithm 1, which optimizes for $H(X\ |\ Y)$ and $H(X)$ simultaneously.
@@
     Update the parameters by gradients: $\phi=\phi+\nabla\mathcal{L}_{1}$
@@
     Update the parameters by gradients: $\phi^{\prime}=\phi^{\prime}+\nabla\mathcal{L}_{2}$
@@
Algorithm 1 can be generalized to any normalizing flows with inner autoregressive structure between $X$ and $Y$. Compared with general autoregressive flows which usually model the autoregressive functions as conditioner neural networks, BNAF has not only the superior expressive power, but also the easy computation of Jacobian matrix and the straightforward deactivation operation given by the block-wise matrix form of autoregressive functions. The theoretical justification based on universal approximation results for Block Autoregressive Flows is provided in an appendix.
4 Numerical Experiments
We implemented several experimental settings from prior work (Belghazi et al., 2018; Song & Ermon, 2019; Poole et al., 2019; Hjelm et al., 2018; Czyż et al., 2023) to evaluate the performance of the proposed estimator. We first focus on the accuracy of the resulting estimates on synthetic Gaussian examples, where the true value of MI can be calculated analytically. In the Appendix C, we report on additional experiments on extremely small-sized dataset and non-Gaussian distributions. The final experiments will be the long-run training behavior on the proposed estimator. All experiments were conduced on a computing cluster using Nvidia Quadro RTX 6000 GPUs.
4.1 MI Estimation on Correlated Multivariate Gaussians
In this experiment, we sampled from two correlated Gaussian random variables $X$ and $Y$, for which the MI can be exactly obtained from their known correlation. The different MI estimators were trained on datasets with varying dimensionality of $X,Y$ ($20$-d, $50$-d and $100$-d), sample size ($32$K, $64$K and $128$K) and true MI to characterize the relative behaviors of every MI estimator. Additionally, we conduct an experiment by applying an element-wise cubic transformation on $y_{i}\rightarrow y_{i}^{3}$. This generates the non-linear dependencies in data without changing the ground truth of MI. The performance of trained estimators are evaluated on a different testing set of $10240$ samples. Czyż et al. (Czyż et al., 2023) mentioned that Gaussians with sparse interactions between $X$ and $Y$ could be a challenging benchmark for MI estimations. We then sample from Gaussians with $\textnormal{Cor}(X1,Y1),\textnormal{Cor}(X2,Y2)>0$ and there is no correlation between any other (distinct) variables. We named it Sparse Gaussian as the covariance matrix $Cov(X,Y)$ is sparse matricx in this case. We assessed our methods along with the following baselines: 1. DEMI (Liao et al., 2020), with the parameter $\alpha=0.5$. 2. SMILE (Song & Ermon, 2019), with three clipping parameters $\tau\in\{5.0,\infty\}$. For $\tau=\infty$, it is equivalent to the MINE Belghazi et al. (2018); 3. InfoNCE (van den Oord et al., 2018), which is the method in contrastive predictive coding (CPC); 4. NWJ (Nguyen et al., 2010), which is the method that based on estimating the likelihood ratios by convex risk minimization; 5. DoE (McAllester & Stratos, 2018), the DoE method, where the distributions is parameterized by isotropic Gaussian (correct) or logistic (misspecified), with three parameters $\tau=1.0$ that clips the gradient norm in training; 6. BNAF, approximating the entropies respectively in the MI using two separate Block Neural Autoregressive Flows; 7. NDoE, BNAF, the proposed method with BNAF structure; 8. NDoE, Real NVP, the proposed method with Real NVP structure. We noticed that the comparison between our method, as a generative model, and other discriminative methods can be difficult since the neural network structure and the model parametrizations are different. To make the comparison as fair as possible, we used the same neural network architecture for all discriminative methods, which is a multi-layer perceptron with an initial concatenation layer, two fully connected layers with $512$ hidden units for each layer and ReLU activations, and a linear layer with a single output. In terms of our proposed method, we constructed the flow with 2 BNAF transformation layers and tanh activations, and a linear BNAF layer to reset the dimensionality. The BNAF layers use $20\times 20$-d, $10\times 50$-d, $6\times 100$-d hidden dimensions for $20$-d, $50$-d and $100$-d data respectively, which is roughly the same as the 512 hidden units in discriminative methods. For Real NVP layers, we let each of the scale and translation functions to be two layers multi-layer perceptron with $128$ hidden units for each layer and ReLU activations. Each MI estimator was trained for $50$ epochs with a mini-batch of $128$. Due to the vanishing and exploding gradient issues in Real NVP, we applied the Adamax optimizer with a fine-tuned learning rate. For all other optimizations, the Adam optimizer with a learning rate of 0.0005 was used. All results were computed over 10 runs on the testing sets generated with different random seeds to ensure robustness and generalizability.
Among all the methods, our proposed model achieved better performance across different dimensionalities and sample sizes. While DoE methods performed well for Gaussian variables, they exhibited a large bias when applied to cubic Gaussians. When comparing NDoE, BNAF with BNAF, we observed that for smaller sample sizes (or insufficient training steps), BNAF exhibits a slight bias across all true MI values. Additionally, for cubic cases, BNAF shows a larger bias when MI is close to zero, an issue not observed with NDoE, BNAF. The work by Song & Ermon (2019) attributed this as a shortcoming of generative models, but we believe our proposed method mitigates this issue, as the bias in entropy estimation vanishes by approximating entropies using the same neural network.
In the cubic cases, NDoE, Real NVP exhibits a larger bias, though it still outperforms discriminative methods, particularly when the sample size is sufficiently large. In the 20-dimensional Gaussian case, SMILE occasionally overestimated MI, which we will further analyze in the long-run training experiments. For the Sparse Gaussian case, both NDoE, BNAF showed small biases when the true MI is small, and BNAF outperformed NDoE, BNAF for larger MI. Both methods consistently outperformed other approaches across different sample sizes. However, NDoE, Real NVP failed to achieve realistic results in the Sparse Gaussian case.

### v3

@@ Abstract
Estimating Mutual Information (MI), a key measure of dependence of random quantities without specific modeling assumptions, is a challenging problem in high dimensions. We propose a novel mutual information estimator based on parametrizing conditional densities using normalizing flows, a deep generative model that has gained popularity in recent years. This estimator leverages a block autoregressive structure to achieve improved bias-variance trade-offs on standard benchmark tasks.
[…]
2983
@@ 1 Introduction
Mutual Information (MI), a measure of dependence of random variables $X$ and $Y$, plays an important role in information theory [6], statistics and machine learning [32, 28, 34, 5], and biology and medical sciences [35, 30]. For random variables $X$ and $Y$ with joint density $p$, the mutual information is defined as
@@
Formally, an MI estimator $\hat{I}_{N}$ depends on independent and identically distributed (i.i.d.) random sample pairs $(X_{1},Y_{1}),\dots,(X_{N},Y_{N})$, and should ideally be unbiased, consistent, and efficient. In addition, such an estimator should be effectively computable from large and high-dimensional data. We propose an unbiased and consistent mutual information estimator based on the difference-of-entropies (DoE) estimator suggested in McAllester and Stratos [23]. This characterization expresses the mutual information as the difference between the entropy of $X$ and the conditional entropy of $X$ given $Y$,
I(X;Y)=H(X)H(X|Y).I(X;Y)=H(X)-H(X\ |\ Y).
@@ 1.1 Overview of previous work
Traditional MI estimators are nonparametric estimators that depend on density estimation and Monte Carlo integration or on the calculation of $k$ nearest neighbors (kNN). Examples include the widely used KSG estimator by Kraskov et al. [18], the nonparametric kNN estimator (kpN) by Lombardi and Pant [22], and improvements of the KSG estimator and a geometric kNN estimator by Gao et al. [12]. These nonparametric methods are fast and accurate for low-dimensional and small-sized problems and are easy to implement. However, they suffer from the curse of dimensionality and do not scale well in machine learning problems since data sets can be relatively large and high-dimensional [27].
@@
More recent parametric methods take advantage of deep learning architectures to approximate variational bounds on MI. These have been categorized into discriminative and generative approaches in Song and Ermon [31]. Some state-of-the-art discriminative approaches include InfoNCE [33], MINE [2], SMILE [31], CCMI [24] and DEMI [20]. van den Oord et al. [33] proposed a contrastive predictive coding (CPC) method that relies on maximizing a lower bound on mutual information. The lower bound involves function approximators implemented with neural networks and is constrained by batch size $N$, leading to a method that is more biased, but with less variance. MINE, on the other hand, is based on the Donsker-Varadhan (DV) lower bound for the KL divergence. The fundamental limitations on approaches based on variational limits were studied by  Song and Ermon [31] and McAllester and Stratos [23], the latter being the motivation for our approach.
@@
Instead of constructing mutual information estimators based on variational lower bounds, Liao et al. [20] proposed a classifier-based estimator called DEMI, where a parametrized classifier is trained to distinguish between the joint density $p(x,y)$ and the product $p(x)p(y)$. Mukherjee et al. [24] proposed another classifier-based (conditional) MI estimator that is asymptotically equivalent to DEMI. However, it still relies on variational lower bounds and is prone to higher error than DEMI for finite samples, as summarized by Liao et al. [20].
@@
Compared with discriminative approaches, generative approaches are less explored in MI estimation problems. A naïve approach using generative models for estimating MI is to learn the entropies $H(X),H(Y)$ and $H(X,Y)$ with three individual generative models, such as VAE [14] or Normalizing Flows [16], from samples. A method for estimating entropy using normalizing flows was introduced by Ao and Li [1]. Estimators based on the individual entropies will be highly biased and computationally expensive since the entropies are trained separately, while it is revealed that considering the enhancement of correlation between entropies in constructing MI estimators can improve the bias [12]. Duong and Nguyen [10] proposes the Diffeomorphic Information Neural Estimator (DINE), that takes advantage of the invariance of conditional mutual information under diffeomorphisms. An alternative approach to estimating MI using normalizing flows was recently proposed by Butakov et al. [4]. This approach takes advantage of the invariance of the point-wise mutual information under diffeomorphisms. In addition, the methods from Butakov et al. [4] allow for the estimation of mutual information using direct, closed-form expressions. Finally, we would like to point to the recently introduced MINDE estimator [11], which is based on diffusion models and represents a complementary approach.
@@
From a practical point of view, the performance of MI estimators is often measured using standard data sets based on Gaussian distributions for which the ground truth is known. Recently, Czyż et al. [7] proposed a collection of benchmarks to evaluate the performance of different MI estimators in more challenging environments.
@@ 2.1 Mutual Information and Entropy
Given a pair of random variables $(X,Y)$, the conditional entropy of $X$ with respect to $Y$ is defined as H(X|Y)=H(X,Y)H(Y)H(X|Y)=H(X,Y)-H(Y), where $H(X,Y)$ is the joint entropy of $(X,Y)$ (not to be confused with the cross-entropy, introduced below). The mutual information can be expressed in terms of entropies via the 3H principle:
@@
The characterization (*) is the basis of the difference-of-entropies (DoE) estimator introduced by McAllester and Stratos [23].
@@
$Q(p,q)\mathrel{\mathop{\ordinarycolon}}=-\mathbb{E}_{p}[\log q(X)].$
One easily checks that the cross-entropy, entropy, and KL divergence are related via
@@
Let $(X,Y)$ be a pair of random variables with joint density $p$. Then
@@
Given data $\{(x_{i},y_{i})\}_{i=1}^{N}$, the resulting difference-of-entropies (DoE) estimator, as suggested by McAllester and Stratos [23], consists of minimizing the objectives
$\begin{split}\hat{Q}(p_{X},q_{X})&=-\frac{1}{N}\sum_{i=1}^{N}\log q_{X}(x_{i}),\\ \hat{Q}(p,q_{X|Y})&=-\frac{1}{N}\sum_{i=1}^{N}\log q_{X|Y}(x_{i}|y_{i})\end{split}$ (4)
@@ 2.2 Normalizing flows
$\begin{split}p_{X}(x)=p_{Z}(g^{-1}(x))\cdot\mathinner{\!\left\lvert\det\textnormal{d}g(g^{-1}(x))\right\rvert}^{-1},\end{split}$
@@
It is known that any continuous distribution with density $p_{X}$ satisfying some mild conditions can be generated from the uniform distribution on a cube $0,1]^{n}$ (and hence, by invertibility, from any other distribution satisfying the same conditions) if the transformation $f$ can have arbitrary complexity [3]. However, as is common with universal approximation results, this result does not translate into a practical recipe [21]. A more practical approach is to use a composition of simple functions implemented by neural networks, which have sufficient expressive power. An obvious but important property of diffeomorphisms is that they are composable. Specifically, let $g_{1},g_{2},\dots,g_{K}$ be a set of $K$ diffeomorphisms and denote by $g=g_{K}\circ g_{K-1}\circ\cdots\circ g_{1}$ the composition of these functions. The determinant of the Jacobian is then given by
@@
In light of equation 2, minimizing the forward KL divergence is equivalent to minimizing the cross-entropy:
$\begin{split}\mathcal{L}(\Theta)&\mathrel{\mathop{\ordinarycolon}}=Q(p_{X}^{*},p_{X}(\cdot;\Theta))=-\mathbb{E}_{p_{X}^{*}}[\log p_{X}(X;\Theta)]\\ &=-\mathbb{E}_{p_{X}^{*}}[\log p_{Z}(f_{\phi}(X);\psi)+\log|\det\textnormal{d}f_{\phi}(X)|].\end{split}$ (6)
@@
$\hat{\mathcal{L}}(\Theta)\mathrel{\mathop{\ordinarycolon}}=-\frac{1}{N}\sum_{j=1}^{N}(\log p_{Z}(f_{\phi}(x_{j});\psi)+\log|\det\textnormal{d}f_{\phi}(x_{j})|).$ (7)
@@ 2.2.1 Block autoregressive flows
Autoregressive flows [15] are normalizing flows with the convenient property that their Jacobian is triangular. Block neural autoregressive flows (B-NAF), introduced by De Cao et al. [8], are flows that are autoregressive and monotone, but that are implemented using a single neural network architecture, rather than relying on conditioner networks. More specifically, a block autoregressive flow is given as a sequence of transformations
@@
where each $B_{ij}^{(k)}\in R^{a_{k+1}\times a_{k}}$ and $g(x)=\exp(x)$ is applied component-wise, to ensure that the entries are positive. We set $a_{0}=a_{\ell+1}=1$. It is not hard to see that the $i$-component of $f(x)$ only depends on $x_{1},\dots,x_{i}$. Since the product of block diagonal matrices with blocks of size $a\times b$ and $b\times c$, respectively, is block diagonal with size $a\times c$, the composition $f$ has a lower triangular Jacobian with positive diagonal entries, and hence is invertible. The determinant of the triangular Jacobian matrix is the product of the diagonal entries $\partial f_{i}/\partial x_{i}$, each of which can be computed as product
@@ 3 Joint estimation of Mutual Information
The Jacobian determinant is the product of the partial derivatives $\partial f_{1}/\partial y$ and $\partial f_{2}/\partial x$ (see Section 2.2.1). Suppose $p(x,y)$ is a standard Gaussian density, so that $p(x,y)=p_{X}(x)p_{Y}(y)$, and that we have data $(x_{i},y_{i})$ from an unknown distribution $q$. The cost function equation 7 for learning a normalizing flow takes the form
$\begin{split}\hat{\mathcal{L}}(\Theta)=&-\frac{1}{N}\sum_{i=1}^{N}\left(\log p_{X}(f_{2}(y_{i},x_{i}))+\log\frac{\partial f_{2}}{\partial x}(y_{i},x_{i})\right)\\ &+\left(\log p_{Y}(f_{1}(y_{i}))+\log\frac{\partial f_{1}}{\partial y}(y_{i})\right).\end{split}$ (8)
The components $f_{1}$ and $f_{2}$ depend on a distinct set of weights in the neural network. Optimizing only the part of equation 8 involving $f_{1}$ on data $\{y_{i}\}$ gives an estimate for the entropy of $Y$, while optimizing the part with $f_{2}$ on data $\{(x_{i},y_{i})\}$ gives rise to an estimate of the conditional entropy H(X|Y)H(X\ |\ Y). Moreover, if we deactivate the weights in off-diagonal blocks (the dashed lines), then optimizing this part on data $\{x_{i}\}$ gives an estimate of $H(X)$. Note that training for H(X|Y)H(X\ |\ Y) and setting the off-diagonal weights to zero does not automatically give an estimator for $H(X)$. It is, however, conceivable that one can begin with a network that approximates $H(X)$ and then optimize the off-diagonal weights to obtain an approximation of $H(X|Y)$.
@@
where now we simply write $p$ for the density of a Gaussian. Optimizing this function gives an estimate of the conditional entropy H(X|Y)H(X\ |\ Y). If, on the other hand, we set the off-diagonal weights to zero and optimize the resulting function $\mathcal{L}_{2}$, we get an estimator for the entropy $H(X)$. This motivates Algorithm 1, which optimizes for H(X|Y)H(X\ |\ Y) and $H(X)$ simultaneously.
@@
  Update the parameters by gradients: $\phi=\phi-\nabla\mathcal{L}_{1}$
@@
  Update the parameters by gradients: $\phi^{\prime}=\phi^{\prime}-\nabla\mathcal{L}_{2}$
@@
Algorithm 1 can be generalized to any normalizing flows with inner autoregressive structure between $X$ and $Y$. Compared with general autoregressive flows which usually model the autoregressive functions as conditioner neural networks, BNAF has not only the superior expressive power, but also the easy computation of the Jacobian matrix and the straightforward deactivation operation given by the block-wise matrix form of autoregressive functions. The theoretical justification based on universal approximation results for Block Autoregressive Flows is provided in Section B of the supplementary materials [26].
@@
4 Numerical Experiments
We implemented several experimental settings from prior work [2, 31, 29, 13, 7] to evaluate the performance of the proposed estimator. We first focus on the accuracy of the resulting estimates on synthetic Gaussian examples, where the true value of MI can be calculated analytically. In the Section C of supplementary materials [26], we report on additional experiments on extremely small-sized datasets and non-Gaussian distributions. The final experiments are the long-run training behavior on the proposed estimator. All experiments were conducted on a computing cluster using Nvidia Quadro RTX 6000 GPUs. The implementation is available at: https://github.warwick.ac.uk/u1774790/nfmi.
4.1 MI Estimation on Correlated Multivariate Gaussians
In this experiment, we sampled from two correlated Gaussian random variables $X$ and $Y$, for which the MI can be exactly obtained from their known correlation. The different MI estimators were trained on datasets with varying dimensionality of $X,Y$ (2020-d, $50$-d and $100$-d), sample size (3232K, 6464K and 128128K) and true MI to characterize the relative behaviors of every MI estimator. Additionally, we conduct an experiment by applying an element-wise cubic transformation on $y_{i}\rightarrow y_{i}^{3}$. This generates the non-linear dependencies in data without changing the ground truth of MI. The performance of trained estimators are evaluated on a different testing set of $10240$ samples. Czyż et al. [7] mentioned that Gaussians with sparse interactions between $X$ and $Y$ could be a challenging benchmark for MI estimations. We then sample from Gaussians with $\textnormal{Cor}(X1,Y1),\textnormal{Cor}(X2,Y2)>0$ and there is no correlation between any other (distinct) variables. We named it Sparse Gaussian as the covariance matrix $Cov(X,Y)$ is a sparse matrix in this case. We assessed our methods along with the following baselines: 1. DEMI [20], with the parameter $\alpha=0.5$. 2. SMILE [31], with three clipping parameters $\tau\in\{5.0,\infty\}$. For $\tau=\infty$, it is equivalent to the MINE [2]; 3. InfoNCE [33], which is the method in contrastive predictive coding (CPC); 4. NWJ [25], which is the method based on estimating the likelihood ratios by convex risk minimization; 5. DoE [23], the DoE method, where the distributions is parameterized by isotropic Gaussian (correct) or logistic (misspecified), with three parameters $\tau=1.0$ that clips the gradient norm in training; 6. BNAF, approximating the entropies respectively in the MI using two separate Block Neural Autoregressive Flows; 7. NDoE, BNAF, the proposed method with BNAF structure; 8. NDoE, Real NVP, the proposed method with Real NVP structure. The implementation example is provided in Section A of supplementary materials [26]. We noticed that the comparison between our method, as a generative model, and other discriminative methods can be difficult since the neural network structure and the model parametrizations are different. To make the comparison as fair as possible, we used the same neural network architecture for all discriminative methods, which is a multi-layer perceptron with an initial concatenation layer, two fully connected layers with $512$ hidden units for each layer and ReLU activations, and a linear layer with a single output. In terms of our proposed method, we constructed the flow with 2 BNAF transformation layers and tanh activations, and a linear BNAF layer to reset the dimensionality. The BNAF layers use $20\times 20$-d, $10\times 50$-d, $6\times 100$-d hidden dimensions for $20$-d, $50$-d and $100$-d data respectively, which is roughly the same as the 512 hidden units in discriminative methods. For Real NVP layers, we let each of the scale and translation functions be two layers multi-layer perceptron with $128$ hidden units for each layer and ReLU activations. Each MI estimator was trained for $50$ epochs with a mini-batch of $128$. Due to the vanishing and exploding gradient issues in Real NVP, we applied the Adamax optimizer with a fine-tuned learning rate. For all other optimizations, the Adam optimizer with a learning rate of 0.0005 was used. All results were computed over 10 runs on the testing sets generated with different random seeds to ensure robustness and generalizability.
Among all the methods, our proposed model achieved better performance across different dimensionalities and sample sizes. While DoE methods performed well for Gaussian variables, they exhibited a large bias when applied to cubic Gaussians. When comparing NDoE, BNAF with BNAF, we observed that for smaller sample sizes (or insufficient training steps), BNAF exhibits a slight bias across all true MI values. Additionally, for cubic cases, BNAF shows a larger bias when MI is close to zero, an issue not observed with NDoE, BNAF. The work by Song and Ermon [31] attributed this as a shortcoming of generative models, but we believe our proposed method mitigates this issue, as the bias in entropy estimation vanishes by approximating entropies using the same neural network.
In the cubic cases, NDoE, Real NVP exhibits a larger bias, though it still outperforms discriminative methods, particularly when the sample size is sufficiently large. In the 20-dimensional Gaussian case, SMILE occasionally overestimated MI, which we will further analyze in the long-run training experiments. For the Sparse Gaussian case, both NDoE, BNAF showed small biases when the true MI is small, and BNAF outperformed NDoE, BNAF for larger MI. Both methods consistently outperformed other approaches across different sample sizes. However, NDoE, Real NVP failed to achieve realistic results in the Sparse Gaussian case. This may be due to the Real NVP architecture capturing fewer intrinsic dependencies between $X$ and $Y$.